Georgia, Inc. has collected the following data on one of its products. The direct materials quantity variance is: Direct materials standard (2 lbs @ $1/lb) $ 2 per finished unit Total direct materials cost variance—unfavorable $ 17,750 Actual direct materials used 60,000 lbs Actual finished units produced 20,000 units

$17,750 unfavorable.

$20,000 favorable.

$17,750 favorable.

$20,000 unfavorable.

$2,250 favorable.
Answer #1

Answer : $20,000 Unfavorable

Actual Units Prodcued = 20,000

Standard Quanity Used = 20,000*2 =40,000 Units

Direct Material Used = 60,000

Direct Material Quantity Variance = ( AQ-SQ)* Standard Price

= (60,000-40,000)*1 = $20,000 Unfavorable